A 0.400 kg pendulum bob passes through the lowest part of its path at a speed of 3.50 m/s. (a) What is the tension in the pendulum cable at this point if the pendulum is 80.0 cm long? N (b) When the pendulum reaches its highest point, what angle does the cable make with the vertical? ° (c) What is the tension in the pendulum cable when the pendulum reaches its highest point?
Answer #1

(a) the tension in string is

T=w+F_{c}

T=mg+\frac{mv^{2}}{r}

T=0.400*9.8+\frac{0.400*3.50^{2}}{0.80m}

T=10.05N

(b)

(KE + PE) 1 + (KE + PEG) 1 +mgyf zmo? +0 = 0+ mgL (1 - cos (max) 1 mgL (1 - cos (max) 1 - cos @max uz 29L v? - 1- 2gL cos max

\theta _{max}=arccos(1-\frac{3.5^{2}}{2-9.8*0.8})

\theta _{max}=77.4^{0}

(c)

T=mgcos(\theta _{max})

T=0.400*9.81cos(77.4^{0})

T=0.856N